At 100 oC, Keq = 1.5E8 for the reaction:

CO(g) + Cl2(g)  COCl2(g)

Using appropriate approximation, calculate the partial pressure of CO at 100 oC at equilibrium in a chamber that initially contains COCl2 at a pressure of 0.260 bar.

bar

The number of significant digits is set to 2; the tolerance is +/-4%
Answer #1

colg) + Cl2 (9) loca(8), Keq = 1:5X108 or, codz (7) → ) + 012(8), keq = 1 1-5x108 = 6:678109 ) Initial pressure of coch = 0.2